OpenStudy (anonymous):

help please . for f(x)= 3x^2 +5 and g(x)=7x-2 Find (f-g)(x)

OpenStudy (zzr0ck3r):

\((f-g)(x)=f(x)-g(x)=3x^2+5-(7x-2)\)

OpenStudy (anonymous):

would that be the answer or is there a next step to that?